OpenStudy (yacoub1993):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16) A. y = x – 6 B. y = x + 6 C. y = x D. y = x + 14

hartnn (hartnn):

can you find the slope of given line y=x+8 ?

OpenStudy (yacoub1993):

how ?

OpenStudy (anonymous):

Do u know if we want one line can parallel to how is that m should be ?:)

OpenStudy (yttrium):

It's in the form of y = mx + b, where m or slope is equal to the coefficient of x. So what's the slope then?

OpenStudy (yacoub1993):

ohh let me calculate

OpenStudy (anonymous):

If u want one line parallel to anothre they thier m should be : m=m' Just this :)

OpenStudy (yacoub1993):

lol i still didnt get :P

OpenStudy (yttrium):

slope is oftenly just the coefficient of x.

OpenStudy (yacoub1993):

is it A

OpenStudy (yttrium):

Are you sure?

OpenStudy (anonymous):

and if u want that passes through use it : y-y2=m(X-X2) X2=2 y2=16 OK ?!:)

OpenStudy (anonymous):

@Yacoub1993:???

OpenStudy (yttrium):

i think you already now the slope now @yacoub1993 just be reminded that it passes through a pt (2,16) so you can formulate the equation using two-slope form which is: y - y1 = m (x - x1) where y1 = 16 and x1 = 2 So, can you decide which one is the answer now?

OpenStudy (anonymous):

slope of the given line y=x+8 is 1.As it is parallel so the slope of equation which we have to find will also be 1.then by problem, (y-y1)=slope(x-x1) (y-16)=1.(x-2) y-x=-2+16 y-x=14

OpenStudy (anonymous):

so ans is D.